OpenStudy (anonymous):

find the volume of the solid generated by rotating y=(x^2)/4 x=2 y=0 about the y axis I am suppose to use disc method

OpenStudy (anonymous):

I built the following integral \[V=\Pi \int\limits_{0}^{1}2-2\sqrt{y}dy\] but I think that I might have done it wrong

OpenStudy (anonymous):

@NotTim

OpenStudy (nottim):

i have no idea sorry

OpenStudy (anonymous):

@Opcode

OpenStudy (opcode):

@jim_thompson5910 Is your best bet...

OpenStudy (anonymous):

I am his fan I asked him if he had a chance to help me and he signed off without responding

OpenStudy (opcode):

@Hero Maybe?

OpenStudy (hba):

V=pi r^2 h y=x^2/4 x^2=4y x=2sqrt{y} x= and x=2sqrt{y} r=2-2sqrt{y} V=pi*r^2*h Take the integral of pi*(2-2sqrt{y})^2*dy

OpenStudy (anonymous):

. . . OMG THERE IS A TYPO IN MY TEXTBOOK

OpenStudy (anonymous):

where I am taking the formula from does not square the funtion . . . so much time wasted